OpenStudy (skittles_for_life6422):

!!!!!!!!!!!!!!!!!!!!!!!!!WILL GIVE MEDAL AND FAN!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!HELP!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! A playground is shaped like a rectangle with a width 5 times its length. what is a simplified expression for the distance between opposite corners of the playground

OpenStudy (studygurl14):

Hi @skittles_for_life6422

OpenStudy (skittles_for_life6422):

hi

OpenStudy (studygurl14):

Do you need help, or do you want me to check your answer?

OpenStudy (skittles_for_life6422):

I need help

OpenStudy (studygurl14):

First step. Draw a rectangle.

OpenStudy (studygurl14):

stupid draw tool isn't working. hold on

OpenStudy (skittles_for_life6422):



OpenStudy (studygurl14):

Next step. It says the width is 5 times the length. Now, we don't know the width or the length. So, if we use L for length, then the width equals 5L. correct?

OpenStudy (skittles_for_life6422):

yup

OpenStudy (studygurl14):

Label the rectangle accordingly.

OpenStudy (skittles_for_life6422):

ok

OpenStudy (skittles_for_life6422):

@StudyGurl14 are you still there

OpenStudy (studygurl14):

yeah, sorry. openstudy lagging

OpenStudy (skittles_for_life6422):

its ok B)

OpenStudy (studygurl14):

you want to find the diagonal

OpenStudy (studygurl14):

use the pytahgoraean theorem.

OpenStudy (skittles_for_life6422):

what is it again

OpenStudy (studygurl14):

\(\large a^2+b^2=c^2\)

OpenStudy (skittles_for_life6422):

oh now i remember

OpenStudy (studygurl14):

:)

OpenStudy (skittles_for_life6422):

\[ l \sqrt{26}, 5l, 26l^2, 6l\]

OpenStudy (skittles_for_life6422):

ANSWER CHOICES sorry took so long, lagging

OpenStudy (studygurl14):

what did you get?

OpenStudy (skittles_for_life6422):

@StudyGurl14 \[6l^2\]